Consider three polarizing filters oriented as show
Answer #1

By malus law from unpolarized to polarized

{I_1}=\frac{1}{2}*{I_o}

from polarized to polarized

{I_2}=I_1*cos^2(\theta)

{I_2}=\frac{I_0}{2}*cos^2(30^o)

{I_2}=\frac{3}{8}I_0

{I_3}=\frac{3}{8}I_0*cos^2(60^o)

{I_3}=\frac{3}{32}I_0

as

I_3=I_f\Rightarrow \frac{I_f}{I_o}=\frac{3}{32}